# Generalized Kummer surfaces over finite fields

### v3

@@ 2.4. Diedonné modules of abelian varieties
There is a contravariant equivalence of categories $M(-)$ between finite group schemes over $k$ of $p$-power order and left $D_{k}$-modules of finite length [Pink, Theorem 28.3].
The $p^{r}$-torsion $A[p^{r}]$ of an abelian variety $A$ over $k$ is a finite group scheme, and the limit
@@
is called the Diedonné module of $A$. The module $M(A)$ is free over $W(k)$ of rank $2\dim A$, and the cotangent space $T^{*}_{0}(A)$ of $A$ at zero is canonically isomorphic to $M(A)/FM(A)$ [Pink, Proposition 28.4]. We have the following analog of Lemma 2.5.
@@
If $f:B\to A$ is an isogeny then, $M(f)$ is injective, and $M=M(f)^{-1}(M(B))$ is a Diedonné submodule of $M(A)\otimes{\mathbb{Q}}$.
@@
We say that the set $\{\lambda\in{\mathbb{Q}}|M_{\lambda}\neq 0\}$ is the set of slopes of $M$. If $k$ is a finite field, and $M=M(A)$ is the Diedonné module of an abelian variety $A$ over $k$, then by [Ma63, Theorem 4.1] the slopes of $M$ and slopes of $A$ coincide. This result motivates a more general definition. An abelian variety $A$ over a perfect field $k$ is supersingular, if all slopes of $M(A)$ are equal to $1/2$, and $A$ is ordinary, if the set of slopes of $M(A)$ is $\{0,1\}$. The following result is due to Tate, Shioda, Deligne, and Oort.
@@
[Oort, Theorem 4.2] Let $k$ be a perfect field of characteristic $p$, and let $E$ be a supersingular elliptic curve over a finite field ${\mathbb{F}}_{q}\subset k$. Then any supersingular abelian surface $A$ over $k$ is isogenous to $E^{2}$ over ${\bar{k}}$. In particular, $\operatorname{End}^{\circ}(A)$ is a subalgebra of $M(2,{\mathbb{H}}_{p})$.
@@
Let $A$ be an ordinary abelian variety over $k$. Then $M(A)\cong M\oplus M^{*}$, where $M$ is a Diedonné module of slope $1$ and rank $\dim(A)$ over $W(k)$, and $T^{*}_{0}(A)\cong M/pM.$
@@
According to [Pink, Proposition 15.4], the group scheme $A[p^{n}]$ is a direct sum of its étale part $X^{*}_{n}$, and the local part $X_{n}$. Since $A$ is ordinary, the order of $X^{*}_{n}$ is $p^{n\dim A}$; therefore, the order of $X_{n}$ is also $p^{n\dim A}$. It follows that
@@ 2.5. Finite group actions on abelian varieties
Let $G$ be a finite group, and an action of $G$ on an abelian variety $A$ is given by a homomorphism $G\to\operatorname{End}(A)^{*}$. The induced action on $T^{*}_{0}(A)$ gives a homomorphism
@@
Let $g\in G$ be an element of order $r$. If the image of $g$ in $\operatorname{GL}_{d}(k)$ is trivial, then $r$ is a power of $p$.
@@
We may assume that $r\neq p$ is the prime and that the action of $g$ on $T^{*}_{0}(A)$ is trivial. The group scheme corresponding to the Diedonné module $M(A)/FM(A)$ is a subscheme of $\ker(g-1)$; therefore, the eigenvalues of $M(g)-1$ are not $p$-adic units. Since the norm
@@
Let $Q$ be a $p$-Sylow subgroup of $G$. Then $G_{1}$ is normal in $Q$, and $Q/G_{1}$ is annihilated by $p$.
@@
The group $G_{1}$ is normal in $G$; therefore, it is normal in $Q$ as well. The quotient $Q/G_{1}$ is isomorphic to a subgroup of $\operatorname{GL}_{2}({\mathbb{F}}_{q})$, and the $p$-Sylow subgroup of $\operatorname{GL}_{2}({\mathbb{F}}_{q})$ is annihilated by $p$. ∎
@@
If the $p$-th Sylow subgroup $Q$ of $G$ is not annihilated by $p$, then $G_{1}$ is non-trivial.
@@
Let $p=2$, and let $G=\operatorname{SL}_{2}({\mathbb{F}}_{5})$. The $2$-nd Sylow subgroup of $G$ is $Q_{8}$; therefore, the filtration given by subgroups $G_{i}$ is non-trivial.
@@ 3. Rigid actions on abelian varieties
Let $\mathcal{C}$ be an isogeny class of abelian varieties over $k$, and let $G$ be a finite group. Denote by $\mathcal{C}_{G}$ the category of abelian varieties from $\mathcal{C}$ with an action of $G$ that fixes the group law. Recall that such an action on an abelian variety $A$ is given by a homomorphism $G\to\operatorname{End}(A)^{*}$. We define the $\operatorname{Hom}$ group as follows:
@@
Let $\mathcal{D}_{G}$ be the following category. Objects of $\mathcal{D}_{G}$ are group homomorphisms $G\to\operatorname{End}^{\circ}(A)^{*}$, where $A$ in $\mathcal{C}$, and a morphism from $i_{A}:G\to\operatorname{End}^{\circ}(A)^{*}$ to $i_{B}:G\to\operatorname{End}^{\circ}(B)^{*}$ is a $\psi\in\operatorname{Hom}^{\circ}(A,B)$ that fits the diagram:
@@
There is a natural functor from $\mathcal{C}_{G}$ to $\mathcal{D}_{G}$ sending an abelian variety $A$ with a $G$-action to the induced homomorphism $G\to\operatorname{End}^{\circ}(A)^{*}$.
@@
for all $\ell$. Consider the natural inclusion $\iota:\operatorname{End}(A)\to T_{G}$ to the lattice generated by $\operatorname{End}(A)$ and the image of $i_{A}(G)$. For all primes $\ell$ the localization $\iota\otimes{\mathbb{Z}}_{\ell}$ is an isomorphism; this implies that $T_{G}=\operatorname{End}(A)$, and $G\subset\operatorname{End}(A)$. We proved that there exists an abelian variety $A$ in $\mathcal{C}_{G}$, and the diagram $A^{\prime}\to A^{\prime\prime}\leftarrow A$ induces an isomorphism in $\mathcal{D}_{G}$. ∎
@@
If there exists an abelian variety $A$ with a $G$-action over $k$, then there exists an abelian variety $B$ of the same dimension over a finite field endowed with a $G$-action and an injective homomorphism $\operatorname{End}^{\circ}(A)\to\operatorname{End}^{\circ}(B)$ such that the $G$-action is given by the composition
@@
If $A$ is supersingular, then $B$ is supersingular as well.
@@
Let ${\mathbb{F}}_{q}$ be the algebraic closure of ${\mathbb{F}}_{p}$ in $k$. There exists a smooth connected affine variety $U$ over ${\mathbb{F}}_{q}$ such that $k$ is a field of functions on $U$. If we make $U$ smaller, then there exists a smooth abelian scheme $\mathcal{A}$ over $U$ such that $A$ is the general fiber of $\mathcal{A}$. Let $B$ be a special fiber of $\mathcal{A}$ over a finite extension of ${\mathbb{F}}_{q}$. Then there exists a natural injection [CMLP, Section 1.8.4.1]:
@@
The Newton polygon of $A$ lies not higher than the Newton polygon of $B$ by the Grothendieck–Katz specialization theorem [Ka78]; therefore, if $A$ is supersingular, then $B$ is also supersingular. ∎
Proposition 3.3.
The action of $G$ on $A$ is rigid if and only if any non-trivial $g\in G$ has only finite number of fixed points.
@@
If $A$ is simple over $k$, then any faithful action of a finite group on $A$ is rigid.
@@
Let ${\mathbb{H}}$ be a semisimple algebra over ${\mathbb{Q}}$. Denote by $\mathcal{C}_{\mathbb{H}}$ the following category. Objects of $\mathcal{C}_{\mathbb{H}}$ are homomorphisms of ${\mathbb{Q}}$-algebras ${\mathbb{H}}\to\operatorname{End}^{\circ}(A)$, where $A$ is in $\mathcal{C}$, and a morphism from $h_{A}:{\mathbb{H}}\to\operatorname{End}^{\circ}(A)$ to $h_{B}:{\mathbb{H}}\to\operatorname{End}^{\circ}(B)$ is an element $\psi\in\operatorname{Hom}^{\circ}(A,B)$ that fits in the diagram:
@@
${\mathbb{Q}}[G]=\oplus_{V}{\mathbb{H}}_{V}.$
${\mathbb{Q}}[G]^{\mathrm{rig}}:=\oplus_{V\text{ is WFP}}{\mathbb{H}}_{V}.$
@@
    Let $g$ be a generator of $G$. If $A$ is an abelian variety with a rigid action of $G$, then all primitive roots of unity of order $n$ are eigenvalues of the action of $g$ on $M(A)$ and on $V_{\ell}(A)$ for all $\ell\neq p$.
@@
According to Theorem 3.5, if the action of $G$ on $A$ is rigid then, ${\mathbb{Q}}(\zeta_{n})$ is a subalgebra of $\operatorname{End}^{\circ}(A)$. By [Ri76, Theorem 2.1.1], $V_{\ell}(A)$ is a free module over ${\mathbb{Q}}(\zeta_{n})\otimes{\mathbb{Q}}_{\ell}$, where $\zeta_{n}$ acts $g$. Hence, $V_{\ell}(A)\otimes\bar{\mathbb{Q}}_{\ell}$ is a free module over
@@
where the sum is over all immersions $\varphi_{\zeta}:{\mathbb{Q}}(\zeta_{n})\to\bar{\mathbb{Q}}_{\ell}$. Any such immersion is uniquely determined by the primitive root of unity of order $n$: $\zeta=\varphi_{\zeta}(\zeta_{n})$. It follows that,
@@
There exists an abelian variety with a rigid action of $G$ in the isogeny class of an abelian variety $A$, if and only if there exists a homomorphism of ${\mathbb{Q}}$-algebras ${\mathbb{Q}}[G]^{\mathrm{rig}}\to\operatorname{End}^{\circ}(A).$
Corollary 3.10.
If there exists a supersingular abelian surface $A$ with a rigid action of $G$ over a perfect field $k$, then there exists a supersingular abelian surface with a rigid action of $G$ over ${\mathbb{F}}_{p^{2}}$. In particular, $G/G_{1}$ is a subgroup of $\operatorname{GL}_{2}({\mathbb{F}}_{p^{2}})$.
@@
The binary dihedral group $Q_{4n}$ of order $4n$ is a group generated by two elements $a$ and $b$ with the following relations:
@@
The only summand corresponding to a faithful representation is ${\mathbb{H}}(n,1)$. This representation is rigid because $b^{2}$ acts as $-1$, and $a$ acts as $\zeta_{2n}$. Finally, ${\mathbb{H}}(2,1)\cong{\mathbb{H}}_{2}$, ${\mathbb{H}}(3,1)\cong{\mathbb{H}}_{3}$, and if $n>3$, then
@@
[Bu] Let $G$ be a group with a representation without fixed points. Then all odd Sylow subgroups of $G$ are cyclic, and the even Sylow subgroup is either cyclic, or isomorphic to $Q_{2^{r}}$.
We say that a cyclic group of order $n$ is small, if $n\in\{2,3,4,5,6,8,10,12\}$.
@@
Let $A$ be an abelian surface with a rigid action by a finite group $G$. Then the order of $G$ divides $240$, and any cyclic subgroup of $G$ is small.
@@
A cyclic subgroup of $G$ generates a commutative subalgebra of $\operatorname{End}^{\circ}(A)$ of dimension not greater than $4$; therefore, any cyclic subgroup is small. Moreover, since for an odd prime $p$ the $p$-Sylow subgroup $G^{(p)}$ is cyclic, we have $p\leq 5$, and $G^{(p)}\cong C_{p}$. We showed that the order of $G$ is $2^{r}m$, where $m$ divides $15$. If the $2$-Sylow subgroup is commutative, then, as before, $r\leq 3$. Otherwise, the subalgebra generated by the $2$-Sylow subgroup of $G$ contains ${\mathbb{Q}}[Q_{2^{r}}]^{\mathrm{rig}}$ as a direct summand. According to Lemma 3.11, $r\leq 4$. ∎

### v4

@@ 2.4. Diedonné modules of abelian varieties
There is a contravariant equivalence of categories M()M(-) between finite group schemes over $k$ of $p$-power order and left $D_{k}$-modules of finite length [Pink, Theorem 28.3].
The $p^{r}$-torsion $A[p^{r}]$ of an abelian variety AA over $k$ is a finite group scheme, and the limit
@@
is called the Diedonné module of AA. The module $M(A)$ is free over $W(k)$ of rank $2\dim A$, and the cotangent space $T^{*}_{0}(A)$ of AA at zero is canonically isomorphic to $M(A)/FM(A)$ [Pink, Proposition 28.4]. We have the following analog of Lemma 2.5.
@@
If $f:B\to A$ is an isogeny, then $M(f)$ is injective, and $M=M(f)^{-1}(M(B))$ is a Diedonné submodule of $M(A)\otimes{\mathbb{Q}}$.
@@
We say that the set $\{\lambda\in{\mathbb{Q}}|M_{\lambda}\neq 0\}$ is the set of slopes of $M$. If $k$ is a finite field and $M=M(A)$ is the Diedonné module of an abelian variety AA over $k$, then by [Ma63, Theorem 4.1] the slopes of $M$ and slopes of AA coincide. This result motivates a more general definition. An abelian variety AA over a perfect field $k$ is supersingular, if all slopes of $M(A)$ are equal to $1/2$, and AA is ordinary, if the set of slopes of $M(A)$ is $\{0,1\}$. The following result is due to Tate, Shioda, Deligne, and Oort.
@@
[Oort, Theorem 4.2] Let $k$ be a perfect field of characteristic $p$, and let $E$ be a supersingular elliptic curve over a finite field ${\mathbb{F}}_{q}\subset k$. Then any supersingular abelian surface AA over $k$ is isogenous to $E^{2}$ over $\bar{k}$}. In particular, $\operatorname{End}^{\circ}(A)$ is a subalgebra of $M(2,{\mathbb{H}}_{p})$.
@@
Let AA be an ordinary abelian variety over $k$. Then $M(A)\cong M\oplus M^{*}$, where $M$ is a Diedonné module of slope 11 and rank $\dim(A)$ over $W(k)$, and $T^{*}_{0}(A)\cong M/pM.$
@@
According to [Pink, Proposition 15.4], the group scheme $A[p^{n}]$ is a direct sum of its étale part $X^{*}_{n}$ and the local part $X_{n}$. Since AA is ordinary, the order of $X^{*}_{n}$ is $p^{n\dim A}$; therefore, the order of $X_{n}$ is also $p^{n\dim A}$. It follows that
@@ 2.5. Finite group actions on abelian varieties
Let $G$ be a finite group, and an action of $G$ on an abelian variety AA is given by a homomorphism $G\to\operatorname{End}(A)^{*}$. The induced action on $T^{*}_{0}(A)$ gives a homomorphism
@@
Let $g\in G$ be an element of order $r$. If the image of $g$ in $\operatorname{GL}_{d}(k)$ is trivial, then $r$ is a power of pp.
@@
We may assume that $r\neq p$ is the prime and that the action of $g$ on $T^{*}_{0}(A)$ is trivial. The group scheme corresponding to the Diedonné module $M(A)/FM(A)$ is a subscheme of $\ker(g-1)$; therefore, the eigenvalues of M(g)1M(g)-1 are not $p$-adic units. Since the norm
@@
Let $Q$ be a $p$-Sylow subgroup of $G$. Then $G_{1}$ is normal in $Q$, and $Q/G_{1}$ is annihilated by pp.
@@
The group $G_{1}$ is normal in $G$; therefore, it is normal in $Q$ as well. The quotient $Q/G_{1}$ is isomorphic to a subgroup of $\operatorname{GL}_{2}({\mathbb{F}}_{q})$, and the $p$-Sylow subgroup of $\operatorname{GL}_{2}({\mathbb{F}}_{q})$ is annihilated by pp. ∎
@@
If the $p$-th Sylow subgroup $Q$ of $G$ is not annihilated by $p$, then $G_{1}$ is not trivial.
@@
Let $p=2$, and let $G=\operatorname{SL}_{2}({\mathbb{F}}_{5})$. The 22-nd Sylow subgroup of $G$ is $Q_{8}$; therefore, the filtration given by the subgroups $G_{i}$ is nontrivial.
@@ 3. Rigid actions on abelian varieties
Let $\mathcal{C}$ be an isogeny class of abelian varieties over $k$, and let $G$ be a finite group. Denote by $\mathcal{C}_{G}$ the category of abelian varieties from $\mathcal{C}$ with an action of $G$ that fixes the group law. Recall that such an action on an abelian variety AA is given by a homomorphism $G\to\operatorname{End}(A)^{*}$. We define the $\operatorname{Hom}$ group as follows:
@@
Let $\mathcal{D}_{G}$ be the following category. Objects of $\mathcal{D}_{G}$ are group homomorphisms $G\to\operatorname{End}^{\circ}(A)^{*}$, where AA in $\mathcal{C}$, and a morphism from $i_{A}:G\to\operatorname{End}^{\circ}(A)^{*}$ to $i_{B}:G\to\operatorname{End}^{\circ}(B)^{*}$ is a $\psi\in\operatorname{Hom}^{\circ}(A,B)$ that fits the diagram:
@@
There is a natural functor from $\mathcal{C}_{G}$ to $\mathcal{D}_{G}$ sending an abelian variety AA with a $G$-action to the induced homomorphism $G\to\operatorname{End}^{\circ}(A)^{*}$.
@@
for all $\ell$. Consider the natural inclusion $\iota:\operatorname{End}(A)\to T_{G}$ to the lattice generated by $\operatorname{End}(A)$ and the image of $i_{A}(G)$. For all primes $\ell$ the localization $\iota\otimes{\mathbb{Z}}_{\ell}$ is an isomorphism; this implies that $T_{G}=\operatorname{End}(A)$, and $G\subset\operatorname{End}(A)$. We showed that there exists an abelian variety AA in $\mathcal{C}_{G}$, and the diagram $A^{\prime}\to A^{\prime\prime}\leftarrow A$ induces an isomorphism in $\mathcal{D}_{G}$. ∎
@@
If there exists an abelian variety AA with a $G$-action over $k$, then there exists an abelian variety $B$ of the same dimension over a finite field endowed with a $G$-action and an injective homomorphism $\operatorname{End}^{\circ}(A)\to\operatorname{End}^{\circ}(B)$ such that the $G$-action is given by the composition
@@
If AA is supersingular, then $B$ is also supersingular.
@@
Let ${\mathbb{F}}_{q}$ be the algebraic closure of ${\mathbb{F}}_{p}$ in $k$. There exists a smooth connected affine variety $U$ over ${\mathbb{F}}_{q}$ such that $k$ is a field of functions on $U$. If we make $U$ smaller, then there exists a smooth abelian scheme $\mathcal{A}$ over $U$ such that AA is the general fiber of $\mathcal{A}$. Let $B$ be a special fiber of $\mathcal{A}$ over a finite extension of ${\mathbb{F}}_{q}$. Then there exists a natural injection [CCO14, Section 1.8.4.1]:
@@
The Newton polygon of AA is not higher than the Newton polygon of $B$ according to the Grothendieck–Katz specialization theorem [Ka78]; therefore, if AA is supersingular, then $B$ is also supersingular. ∎
Proposition $3.$
The action of $G$ on AA is rigid if and only if any non-trivial $g\in G$ has only finite number of fixed points.
@@
If AA is simple over $k$, then any faithful action of a finite group on AA is rigid.
@@
Let $\mathbb{H}$} be a semisimple algebra over $\mathbb{Q}$}. Denote by $\mathcal{C}_{\mathbb{H}}$ the following category. Objects of $\mathcal{C}_{\mathbb{H}}$ are homomorphisms of $\mathbb{Q}$}-algebras ${\mathbb{H}}\to\operatorname{End}^{\circ}(A)$, where AA is in $\mathcal{C}$, and a morphism from $h_{A}:{\mathbb{H}}\to\operatorname{End}^{\circ}(A)$ to $h_{B}:{\mathbb{H}}\to\operatorname{End}^{\circ}(B)$ is an element $\psi\in\operatorname{Hom}^{\circ}(A,B)$ that fits the diagram:
@@
    Let $g$ be a generator of $G$. If AA is an abelian variety with a rigid action of $G$, then all primitive roots of unity of order $n$ are eigenvalues of the action of $g$ on $M(A)$ and on $V_{\ell}(A)$ for all $\ell\neq p$.
@@
According to Theorem 3.5, if the action of $G$ on AA is rigid, then ${\mathbb{Q}}(\zeta_{n})$ is a subalgebra of $\operatorname{End}^{\circ}(A)$. By [Ri76, Theorem 2.1.1], $V_{\ell}(A)$ is a free module over ${\mathbb{Q}}(\zeta_{n})\otimes{\mathbb{Q}}_{\ell}$, where $\zeta_{n}$ acts $g$. Hence, $V_{\ell}(A)\otimes\bar{\mathbb{Q}}_{\ell}$ is a free module over
@@
where the sum is over all immersions $\varphi_{\zeta}:{\mathbb{Q}}(\zeta_{n})\to\bar{\mathbb{Q}}_{\ell}$. Any such immersion is uniquely determined by the primitive root of unity of order $n$: $\zeta=\varphi_{\zeta}(\zeta_{n})$. It follows that
@@
If there exists a supersingular abelian surface AA with a rigid action of $G$ over a perfect field $k$, then there exists a supersingular abelian surface with a rigid action of $G$ over ${\mathbb{F}}_{p^{2}}$. In particular, $G/G_{1}$ is a subgroup of $\operatorname{GL}_{2}({\mathbb{F}}_{p^{2}})$.
@@
The binary dihedral group $Q_{4n}$ of order $4n$ is the group generated by two elements $a$ and $b$ with the following relations:
@@
The only summand corresponding to a faithful representation is ${\mathbb{H}}(n,1)$. This representation is rigid because $b^{2}$ acts as 1-1, and $a$ acts as $\zeta_{2n}$. Finally, ${\mathbb{H}}(2,1)\cong{\mathbb{H}}_{2}$, ${\mathbb{H}}(3,1)\cong{\mathbb{H}}_{3}$, and if $n>3$, then
@@
[Bu] Let $G$ be a group with a representation without fixed points. Then all odd Sylow subgroups of $G$ are cyclic, and the even Sylow subgroup is cyclic or isomorphic to $Q_{2^{r}}$.
We say that a cyclic group of order $n$ is small if $n\in\{2,3,4,5,6,8,10,12\}$.
@@
Let AA be an abelian surface with a rigid action by a finite group $G$. Then the order of $G$ divides $240$, and any cyclic subgroup of $G$ is small.
@@
A cyclic subgroup of $G$ generates a commutative subalgebra of $\operatorname{End}^{\circ}(A)$ of dimension not greater than 44; therefore, any cyclic subgroup is small. Moreover, since for an odd prime $p$ the $p$-Sylow subgroup $G^{(p)}$ is cyclic, we have $p\leq 5$, and $G^{(p)}\cong C_{p}$. We showed that the order of $G$ is $2^{r}m$, where $m$ divides $15$. If the 22-Sylow subgroup is commutative, then, as before, $r\leq 3$. Otherwise, the subalgebra generated by the 22-Sylow subgroup of $G$ contains ${\mathbb{Q}}[Q_{2^{r}}]^{\mathrm{rig}}$ as a direct summand. According to Lemma 3.10, $r\leq 4$. ∎